OpenStudy (anonymous):

a charge particle is free to move in an electric field ....what kind of path it wil travel...with respect to the line of force?????

OpenStudy (anonymous):

The force on a charged particle is from an electric field \(F=qE\cos(\theta)\) Lets consider two extreme cases. First lets see what happens if the angle between the electric field and the velocity vector is zero - they both point in the same direction - \(\theta = 0\): \(F=qE\cos(0)=qE\) since \(\cos(0) = 1\). This means that a force equal to \(qE\) is exerted in the same direction it is going - it will accelerate. Now lets see what happens when the particle is moving in the opposite direction to the electric field, or \(\theta = 180\). \(F=qE\cos(0)=-qE\) since \(\cos(180) = -1\). This means that a force equal to \(qE\) is exerted in the opposite direction it is going - it will decelerate. You can see this in this picture